Solve the IVP 3 [31 = [ +5] [i] [OF
Answer #1

The given IvP is 2-57 5-8 ac where x = and A= jie. X=AX with initial Condition -X-6] - [:-] Let X=ę ert (20) be a trial soluThen 2 -5 5 - 9 = -3 22 224-522 = -321 and 534 -8X2=-322 which are equivalent to 74-72=0 . 74-72 If x=0 then s4=c 13- [8] =of test and eat both to be zero. From the term in e 3t we find that =0 Hence there is no solution of the system (1) of the fosn, 5%2=1 > n-n2=/5 Take Mk Then ny = 5 +k > where KER. n= [] - [***] - *[:-[:] x=ętest the 3t 3t + est te 3t +k + est e-3t TInitial condition given: [sco [you]= [-] Cine, come pending Therefore from (4 we get [1] = a[i] + c2[%] G + Ś C2 = 3 CI= -1 :